A party of fishermen rented a boat for $240. Two of the men had to withdraw from the party and, as a result, the share of each of the others was increased by $10. How many fishermen were in the original party?
Answer by macston(5194)

.
x=original number of fishermen
.
Original cost=$240/x
New cost=$240/(x-2)
.
New cost - original cost=$10
($240/(x-2))-$240/x=$10
$240x/(x-2)-$240=$10x
$240x-$240x+$480=$10x^2-$20x
$480=$10x^2-$20
48=x^2-2x
0=x^2-2x-48
0=(x-8)(x+6)
x-8=0 OR x+6=0
x=8 OR x=-6 There cannot be -6 fishermen, so:
ANSWER: The number of fishermen in the original party was 8.